Calculate the number of milliliters of solution required to provide the following:


a. 17.5 g of sulfuric acid from a 2.50 N solution in reactions that replace both hydrogen ions.


b. 24.2 g of calcium chloride from a 1.25 N solution in reactions that replace both chloride ions.


1
Expert's answer
2022-05-16T23:55:02-0400

"c((H_2SO_4)=2.5 n=2.5\/2=1.25 M"

"M(H_2SO_4)=2x1+32+16x4=98 g\/mole"

"n(H_2SO_4)=17.5 \/98=0.179 mole"

"V(H_2SO_4)=0.179\/1.25=0.1428 l=142.8 ml"

b."c(CaCl_2)=1.25 n=1.25\/2=0.625 M"

"M(CaCl_2)=40+2x35.5=111 g\/mole"

"n(CaCl_2)=24.2\/111=0.218 mole"

"V(CaCl_2)=0.218\/0.625=0.3488 l=348.8 ml"